Mass of NiCl2 in 5 mL saturated solution = 2.73 g
Solubility (g/L) of NiCl2 = 546 g/L
Molar solubility of NiCl2 = 4.44 M
[Ni2+] = 4.44 M
[Cl-] = 8.88 M
Ksp for NiCl2 = 350.1

Why is it reasonable that the Ksp of NiCl2 should be so high?
Answer #1

Mass of Nich in 5 ure Saturated solution = 2.73g Solubility = 2.739/5 me = 0.546 g/me -0.5 46x1600) 9/L = 546 g/ 2+ Molar SolNielz + 6420 [Ni (426) 6) eh for complere formation solubility of Nich is very large and ksp is very large.